Circular No. 8405

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ej AND 2004ek Two apparent supernovae have been found on unfiltered CCD images:

2004ej in NGC 3095 by L. A. G. Monard (IAUC 8382); and 2004ek in UGC 724 by T. Boles (IAUC 8401) and by T. Puckett and L. Cox (IAUC 8392).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004ej Sept. 10.134 10^{h}00^{m}02.^{s}84 −31^{o}33′25′′.5 16.1 38′′ W, 15′′ S
2004ek Sept. 9.973 1 09 58.51 +32 22 47.7 17.1 11′′ W, 41′′ N

Magnitudes by Monard for SN 2004ej: Digitized Sky Survey, [20.5 (red); July 12.7 UT, [18.0; Aug. 3.7, [17.0; Sept. 12.133, 15.9. Magnitudes for SN 2004ek (tabulated data above by Boles): 2003 Oct. 7, [20.0 (Puckett); 2004 Aug. 25, [19.5 (Boles); Sept. 11.25, 16.8 (Puckett); 12.20, 16.8 (Puckett).

V1187 SCORPII T. R. Geballe, R. S. Fisher, T. L. Hayward, and S. Chan, Gem-

ini Observatory, report new 7–13-μm spectroscopy of V1187 Sco at the Gemini Northern Telescope on Sept. 7 UT using the low-resolution grating (resolution 200) in the instrument ‘MICHELLE.’ The rich hydrogen recombination-line spectrum has weakened by a factor of ∼ 3 since the previous measurement of Aug. 11 (IAUC 8386). The H I 7-6 line currently has a flux of ∼ 5 × 10^{−16} W/m^{2}. Line widths (FWHM ∼ 4000 km/s) are unchanged since the earlier spectrum. The 12.8-μm line of Ne II has appeared since the earlier spectrum and has a flux that is one-sixth that of H I 7-6. The early appearance of this line indicates that V1187 Sco belongs to the class of ONeMg novae. As the strength of the Ne II line is expected to increase over the next few weeks, further observations are encouraged.

SUPERNOVA 2004eb IN NGC 6387 D. C. Leonard, A. Soderberg, A. Gal-Yam, B. Cenko, D. B. Fox, D.

S. Moon, and D. J. Sand, California Institute of Technology, report that inspection of a preliminarily reduced CCD spectrum (range 550–900 nm, resolution 1.8 nm), obtained as part of the Caltech Core-Collapse Project (cf. http://www.astro.caltech.edu/˜ avishay/cccp.html) with the Palomar 5m telescope on Sept. 9 UT, reveals that SN 2004eb (cf. IAUC 8396) is a supernova, most likely of type II. The spectrum, heavily contaminated by the underlying host galaxy, displays a blue continuum with broad emissiondominated P-Cyg features at the locations of Hα and He I 587.6-nm.

2004 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green